OpenStudy (anonymous):

Rene is going to the lake to visit some friends. If the lake is 60 miles away, and Rene is driving at 40 miles per hour the entire time, how long will it take her to get to the lake?

OpenStudy (anonymous):

Since the four possible answers are in minutes, it would help greatly if you converted the speed into miles per minute. 40 miles/hour * 1hour/60minutes = 40/60 = 2/3 mile/minute since we will use this in a formula, we leave it in fraction form to obtain the highest degree of accuracy. Now that we have the speed in miles/minute, when we calculate the time, it will also be in minutes. time = distance/speed = 60 / (2/3) dividing by a fraction is the same as multiplying by the inverse of that fraction = 60 * 3/2 = 180/2 = 90 minutes

OpenStudy (anonymous):

thanks

OpenStudy (emmamink):

It will take her a hour and a half or 90 minutes.

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

I have the answer now thanks :)